HUBBLE SPACE TELESCOPE DAILY REPORT #2281 PERIOD COVERED: 0000Z (UTC) 01/06/99 - 0000Z (UTC) 01/07/99 Daily Status Report as of 007/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Twelve Sets of NIC/1/2/3 8093 (DARKs: Warming Up - Part II) The Near Infrared Camera (Cameras One, Two and Three) was used to make a series of dark calibration observations. The purpose of this proposal is to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. As described in 2.1 and HSTAR 6866, the acquisition for the next to last iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed STIS/CCD/MA1 7432 (Morphological and Spectral Evolution of Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1) were used to make observations to study the morphological and spectral evolution of the galaxy 3C196.1-1. A library of ultraviolet images of normal galaxies at low redshifts is essential to our understanding of the properties and evolution of the high redshift universe. The observations were completed as planned, and no problems were reported. 1.4 Completed Three Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II parallel archive proposal continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FGS 7678 (Monitor the Long-Term Stability of FGS-1R) The goal of these measurements is to monitor the internal positional stability of the FGSs by measuring many stars in the M35 check field repeatedly over a one year period. The results of these measurements are the stability characteristics of the FGS to allow an estimate of differential corrections to OFAD and Plate Scale. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 7763 (A Test For Products of Oxygen Burning in the Crab Nebula) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the Crab Nebula (a.k.a., M-1 or NGC-1952) to test for products of oxygen burning. Young supernova remnants are excellent laboratories for investigating how stars make elements, and the Crab Nebula can provide unique information about the supernova process and associated heavy element production. A wavelength calibration observations was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Four Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.12 Completed WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, whilst producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study using the WF/PC-2 will remedy this situation by undertaking redder, line-free observations of AGN 1215-033. As described in 2.1 and HSTAR 6866, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the proposal completed with no further reported problems. 1.13 Completed Two Sets of STIS/CCD/MA1/MA2 7303 (UV Spectroscopy of Wolf-Rayet-Type Central Stars in the Magellanic Clouds) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make spectrograph observations of the nebulei and Wolf-Rayet class stars in The Large Magellanic Cloud (LMC-MA17-C and LMC-SM61). Consistent modelling of the nebulae will add further information about the evolutionary past of the central stars. A wavelength calibration image was also taken. The observations were completed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 The acquisition at 006/182318Z defaulted to fine lock backup on FGS-1 only when FGS-3 could not achieve fine lock. HSTAR 6866 was written. The proposals described in 1.2 and 1.12 may have been affected. Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1. Two TTRs were written: one for a data inversion hit at 006/1247Z and another for missing minor frames (20 at 006/1609Z and 50 at 007/0928Z. As predicted, the NICMOS FPA temperatures reached the saturation point of 78.1 degrees K early yesterday morning. HSTAR 6861 was written. Real-time operations personnel sent the commands to inhibit filter wheel operations at approximately 006/1432Z. Memory monitors confirmed that the filter wheels were inhibited. NICMOS will continue execution of the engineering proposals until next week's SMS begins execution at the beginning of day 011. At that time, the NICMOS will be transitioned to the SAA operate mode and remain there until SM3 when the installation of the NICMOS Cooling System is expected. Another period of CCS commanding to the spacecraft was successfully concluded at 006/2250Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH